Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

Sales $ 1,920,000
Variable expenses 960,000
Contribution margin 960,000
Fixed expenses 180,000
Net operating income $ 780,000

4-b. Assume the president expects this year's sales to increase by 19%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $72,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $2.00 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $780,000 net operating income as last year?
Answer #1

ANSWER :-

Degree of operating leverage = contribution margin / net income

= 960,000 / 780,000

= 1.23

or = 1.24

4 (b)

Percentage increase in net income = 19% * 1.24

=23.56

or =23.56%

  

5 (a)

unit sale = 19,20,000 / 80

= 24000

now unit sales =24000*125%

=30000

sales (24000*69.6) =1670400

Variable expense (24000*40) = 960000

Contribution margin = 710400

fixed expense (180,000+72000 )= 252000

Net operating income = 458400